# Knowledge Bridger: Towards Training-Free Missing Modality Completion

### v1

@@ 3.1 Setup
Evaluation Metrics. For multimodal multi-label classification on three datasets, we report the performance of comparison methods using the macro F1-score and mean Average Precision (mAP). We report the average result of five different random seeds. Additionally, we introduce the mean similarity score (denoted SS) to evaluate the generation quality of our method and other imputation-based methods. Specifically, we utilize the vision and text backbones of the pre-trained CLIP [39] to compute embeddings for all ground-truth and generated missing modalities. The cosine similarity between each pair of embeddings is then computed and averaged, with scores ranging from 0 to 100. A higher score indicates better generation quality.
@@
Method F1 mAP SS F1 mAP SS F1 mAP SS F1 mAP SS F1 mAP SS F1 mAP SS
Baseline (complete) F1: 78.3 |||| mAP: 84.6 |||| SS: - F1: 56.2 |||| mAP: 62.7 |||| SS: -
@@
Method F1 mAP SS F1 mAP SS
Baseline (complete) F1: 57.0 |||| mAP: 75.7|||| SS: -
@@ 3.2 Quantitative Analysis
In Tables 1 and 2, we present the results of our method compared to others across different missing data ratios, as measured by F1, mAP, and mean similarity score (SS) on three datasets [^3] More results are presented in the appendix.. To comprehensively evaluate the performance of various approaches, we introduce two baselines: one trained on complete data and another trained on data with missing entries removed. Additionally, to fair comparison regarding the scale of model parameters, the proposed method based on the Qwen-VL-2B were incorporated.
Results from general domain datasets (COCO-2014 and MM-IMDb) indicate that most MMC methods outperform the baseline where missing data is simply removed. Our method demonstrates superior performance across different missing rates, particularly at a missing rate of 0.7, where it shows significant improvements in both F1 and mAP metrics. We attribute these enhancements to the synthetic data generated by our method, which appears to bolster downstream task performance. This phenomenon is also observed in [14, 47]. However, it should be noted that the synthetic data may lead to a slight decline in mean similarity score.
@@ 3.3 Ablation Study
Q3: Does generating more missing data improve performance? We evaluate the impact of different numbers of generation candidates in step 2 of our method on the MM-IMDb and IU X-ray datasets. Specifically, we conduct the settings with generation candidates $n=\{1,5,10,15\}$, and the results are presented in Table A. The findings indicate that generating only one candidate significantly reduces the missing generation quality. While using a higher number of generation candidates can lead to slight improvements across all metrics, it also substantially increases inference time. Therefore, setting $n=5$ provides a balance between inference time and performance.
@@ I.1 Table of Quantitative Results
Method F1 mAP SS
Baseline (complete) F1: 57.0 |||| mAP: 75.7|||| SS: -

### v2

@@ 3.1 Setup
Evaluation Metrics. For multimodal multi-label classification on three datasets, we report the performance of comparison methods using the macro F1-score and Average Precision (AP). We report the average result of five different random seeds. Additionally, we introduce the mean similarity score (denoted SS) to evaluate the generation quality of our method and other imputation-based methods. Specifically, we utilize the vision and text backbones of the pre-trained CLIP [41] to compute embeddings for all ground-truth and generated missing modalities. The cosine similarity between each pair of embeddings is then computed and averaged, with scores ranging from 0 to 100. A higher score indicates better generation quality.
@@
Method F1 AP SS F1 AP SS F1 AP SS F1 AP SS F1 AP SS F1 AP SS
Baseline (complete) F1: 78.3 |||| AP: 84.6 |||| SS: - F1: 56.2 |||| AP: 62.7 |||| SS: -
@@
Method F1 AP SS F1 AP SS
Baseline (complete) F1: 57.0 |||| AP: 75.7|||| SS: -
@@ 3.2 Quantitative Analysis
In Tables 1 and 2, we present the results of our method compared to others across different missing data ratios, as measured by F1, AP, and mean similarity score (SS) on three datasets [^3] More results (more modalities) are presented in the appendix.. To comprehensively evaluate the performance of various approaches, we introduce two baselines: one trained on complete data and another trained on data with missing entries removed. Additionally, to fair comparison regarding the scale of model parameters, the proposed method based on the Qwen-VL-2B were incorporated.
Results from general domain datasets (COCO-2014 and MM-IMDb) indicate that most MMC methods outperform the baseline where missing data is simply removed. Our method demonstrates superior performance across different missing rates, particularly at a missing rate of 0.7, where it shows significant improvements in both F1 and AP metrics. We attribute these enhancements to the synthetic data generated by our method, which appears to bolster downstream task performance. This phenomenon is also observed in [16, 49]. However, it should be noted that the synthetic data may lead to a slight decline in mean similarity score.
@@ 3.3 Ablation Study
Q3: Does generating more missing data improve performance? We evaluate the impact of different numbers of generation candidates in step 2 of our method on the MM-IMDb and IU X-ray datasets. Specifically, we conduct the settings with generation candidates $n=\{1,5,10,15\}$, and the results are presented in Table 4. The findings indicate that generating only one candidate significantly reduces the missing generation quality. While using a higher number of generation candidates can lead to slight improvements across all metrics, it also substantially increases inference time. Therefore, setting $n=5$ provides a balance between inference time and performance.
@@ I.1 Table of Quantitative Results
Method F1 AP SS
Baseline (complete) F1: 57.0 |||| AP: 75.7|||| SS: -
@@ I.2 Table of Ablation Study
I.3 More Modalities Results
The proposed method primarily focuses on image and text modalities to facilitate the evaluation of the proposed method, as these modalities are well-supported by the community and computationally efficient when using large vision-language models. However, our approach imposes no constraints on the modality encoders, allowing it to be easily generalized to other modalities. To validate this, we conducted experiment on a multimodal sarcasm detection dataset [11], which includes audio, vision, and text modalities. Using Unified-IO [10] as the backbone and keeping other configurations unchanged, our method demonstrated effectiveness even when extended to these modalities, as shown in Table  7.
Missing Rate $\eta$ 0.3 0.5
Method F1 AP SS F1 AP SS
Baseline (complete) F1: 62.4 |||| mAP: 64.7|||| SS: -
Baseline (remove missing) 57.1 59.3 - 53.8 54.6 -
DiCMoR (CVPR’23) 55.5 57.6 11.7 51.3 52.9 10.2
Ours (Unified-IO 7B) 58.3 60.1 13.3 54.7 55.8 11.4
Table 7: Quantitative results (%) on sarcasm datasets.